OpenStudy (anonymous):

The slope of a line is 0, and the y-intercept is 6. What is the equation of the line written in slope-intercept form? A) y = 6 B) y = 6x C) y = x + 6

OpenStudy (anonymous):

what is slope intercept form??

OpenStudy (tkhunny):

@cheyenneireland So far, you have not shown any work or knowledge. You cannot pass this class with this level of performance. Please demonstrate something. Show us that you know something about linear equations.

OpenStudy (anonymous):

I don't remember it, i am just in summer school trying to make up credits from freshman year and i am a senior now. i only have two days to finish all of this… its not much time to learn its all again!

OpenStudy (anonymous):

i probably know the formula just not the name. i am really confused

OpenStudy (anonymous):

is it y=mx+b??

OpenStudy (tkhunny):

Well, then it appears you are toast. It isn't magic and we can't cram it into your head. All the previous problems used the Slope-Intercept Form. You should be familiar with it. y = mx + b You are given the Slope = 0, this gives: y = (0)x + b = b You are given the y-intercept = 6, this gives: y = 6 Find the information and substitute. Make another post and YOU solve that one. Don't make anyone else do it. We cannot be there on your exam. You must do it by yourself. Stop making excuses for why you cannot do it. Now is the time!

OpenStudy (anonymous):

y=(o)x+(6) so would it be y=6?

OpenStudy (anonymous):

i already took my exam on this section i am just re doing some of the homework to help my grade. @tkhunny

OpenStudy (tkhunny):

You are not addressing the issue. You are NOT doing the homework. Do one from start to finish with no input from anyone. THEN you will be doing the homework.